### v1

# Out-of-Distribution Generalization in Time Series: A Survey

Abstract
Time series frequently manifest distribution shifts, diverse latent features, and non-stationary learning dynamics, particularly in open and evolving environments. These characteristics pose significant challenges for out-of-distribution (OOD) generalization. While substantial progress has been made, a systematic synthesis of advancements remains lacking. To address this gap, we present the first comprehensive review of OOD generalization methodologies for time series, organized to delineate the field’s evolutionary trajectory and contemporary research landscape. We organize our analysis across three foundational dimensions: data distribution, representation learning, and OOD evaluation. For each dimension, we present several popular algorithms in detail. Furthermore, we highlight key application scenarios, emphasizing their real-world impact. Finally, we identify persistent challenges and propose future research directions. A detailed summary of the methods reviewed for the generalization of OOD in time series can be accessed at https://tsood-generalization.com.
Index Terms:
Time series, out-of-distribution generalization, deep learning.
I Introduction
Time series are widely applied in many real-world scenarios, such as financial analysis [1, 2] autonomous driving [3, 4], healthcare [5, 6], and industrial fault diagnosis [7, 8]. The key to most time series methods lies in extracting useful features. These methods are typically trained under the assumption of independent and identically distributed (IID) data, meaning that the training data (observed) and test data (unobserved) are assumed to follow the same independent distribution [9]. However, this assumption does not always hold in practice. Moreover, collecting data from all possible distributions is both time-consuming and labor-intensive. It is often difficult to achieve due to unknown latent factors or data scarcity. Therefore, enhancing the generalization ability of time series methods under out-of-distribution (OOD) scenarios becomes particularly important.
Figure 1: An example of the TS-OOG problem. The test set includes some data with a distribution similar to the training set and unseen time series distributions. If the model fails to account for distribution shifts, its performance on OOD data significantly degrades (top-right plot). After improving generalization, performance is enhanced (bottom-right plot). Therefore, it is crucial to have the capability to adapt to unseen data distributions.
With the rapid advance of deep learning technologies, out-of-distribution generalization in time series (TS-OOG) has garnered increasing attention [10, 11]. As illustrated in Fig. 1, the objective of TS-OOG is to train a learning model on data from a specific period (or historical stage) so that it can effectively generalize to unseen time series distributions. For example, given a set of financial market data from previous years, TS-OOG requires that the trained model be capable of accurately predicting future market trends, even if the distribution of future data shifts significantly due to changes in the economic environment, policy adjustments, or technological advances [12, 13, 14]. Over the past few years, TS-OOG has achieved remarkable progress in various fields. However, there remains a dearth of comprehensive surveys that systematically introduce its core concepts, method classifications, evaluation, and the existing challenges.
Figure 2: The research roadmap for TS-OOG categorizes representative methods in chronological order. For a more comprehensive summary, please refer to the main text.
In this paper, we present the first survey on TS-OOG, highlighting recent advances with a particular focus on data distributions, representation learning, evaluation methods, application scenarios, and future research directions. We organize representative TS-OOG methods in chronological order, as illustrated in Fig. 2. Since 2020, research methodologies have continuously diversified. Over the past two years, large time-series models have become a focal point. By offering a thorough overview, we hope that support researchers and inspire further studies in this field and related areas.
Related Surveys. Our survey differs significantly from the existing literature in scope and technical focus. Liu et al. [15] provide an overview of OOD generalization, emphasizing causal mechanisms and stable architectures. Concurrently, Li et al. [16] focus on graph-structured data using geometric learning paradigms, whereas our research introduces a systematic framework for TS-OOG. This domain poses unique challenges like non-stationary distributions, and dynamic feature interactions. Yang et al [17] explored the challenges of OOD in the era of large-scale language modeling. Crucially, we introduce a novel three-dimensional taxonomy, establishing foundations distinct from previous approaches.
Our Contributions. This survey provides a comprehensive review of TS-OOG, with the following specific contributions:
  • To systematically and comprehensively summarize existing research on TS-OOG, we propose an innovative classification framework that categorizes related studies across three key dimensions.
  • We substantially expand and refine the research landscape by introducing several important new categories, including decoupling-based methods, invariant-based methods, ensemble-based learning, and large time-series models, as illustrated in Fig. 3. Additionally, we incorporate a wide range of recent studies (over 50% of new work).
  • We conduct an in-depth analysis of the foundations, techniques, and real-world applications of various TS-OOG approaches, providing a clear understanding of their intrinsic relationships and distinctions. And we have established an open-source code repository [^1] https://tsood-generalization.com, to support further research and practical applications in this field.
Figure 3: A comprehensive taxonomy of TS-OOG methods.
Organization. The remainder of this paper is organized as follows: In Section II, we introduce the background of TS-OOG. In Section III, we present the taxonomy methods for TS-OOG, covering data distribution, representation learning, and evaluation frameworks. Section IV highlights several typical application scenarios. Section V discusses the challenges and future directions. Finally, in Section VI, we summarize the survey.
TABLE I: Notations used in this survey.
Notation Description Notation Description
$X_{t}$ Observations at time point $t$ $y$ Label
$Y_{t}$ Predictions at time point $t$ $\mathcal{Y}$ Label space
$\mathcal{X}$ Input space $\mathcal{L}$ Loss function
$\pi_{i}$ The weight of the $i$-th domain $\theta$ Parameter
$\mathcal{D}$ Domain $f(\cdot)$ Predictive model
$\mathbb{P}(\cdot)$ Distribution $\epsilon$ Error
II BACKGROUND
II-A Definition of Time Series Data
Time series is a sequence of data points ordered by time, typically representing measurements or observations taken at successive, equally spaced points in time. Mathematically, a time series $X=\{X_{t}\}_{t=1}^{T}$ consists of $T$ observations where each $X_{t}\in\mathbb{R}^{p}$ represents the value of the time series at time $t$, and $p$ denotes the dimensionality of each observation (i.e., the number of features at each time step). Data augmentation operations such as the flipping operation generate a new sequence $X_{t}^{\prime}=-X_{t}$. For cases [18, 19, 20], the time series $X_{t}$ is decomposed as $X_{t}=\tau_{t}+s_{t}+r_{t}$, where $\tau_{t}$ represents the trend component, $s_{t}$ denotes the seasonal component, and $r_{t}$ signifies the residual component.
Consider a time-series training dataset $\mathcal{D}^{\text{train}}=\{(X_{i},Y_{i})\}_{i=1}^{T}$, where each $X_{i}\in\mathcal{X}\subset\mathbb{R}^{p}$ is a $p$-dimensional instance, and $Y_{i}\in Y=\{1,\cdots,C\}$ is its label. Assume the dataset comprises $N$ latent domains, where $1\leq N\leq T$. Each latent domain $i$ has an associated joint distribution $\mathbb{P}^{i}(X,Y)$ over $\mathcal{X}\times\mathcal{Y}$. The weights of these distributions are denoted by $\pi_{i}$, satisfying $\sum_{i=1}^{N}\pi_{i}=1$ and $\pi_{i}>0$. The overall distribution of the time series is expressed as a weighted sum of the distributions across all latent domains:
$\mathbb{P}^{\text{train}}(X,Y)=\sum_{i=1}^{N}\pi_{i}\mathbb{P}^{i}(X,Y).$ (1)
Here, $\mathbb{P}^{\text{train}}(X,Y)$ aligns with the joint distribution of the dataset over the input-output space $\mathcal{X}\times\mathcal{Y}$. Each latent domain is represented by a set of samples following its distribution. These samples can be freely interchanged within their respective domains.
We provide a list of commonly used notations in Table I.
II-B The Formalization of OOD Generalization
II-B1 General Domains
OOD generalization is considered a branch of transfer learning, aiming to train models that can adapt to unseen targets or domains, even when their distributions differ from the training data distribution [21, 22]. The core objective is to learn a generalized model from the training dataset $\mathcal{D}^{\text{train}}$ that can maintain strong predictive performance on an inaccessible target dataset $\mathcal{D}^{\text{test}}$ [9, 23, 24]. The input and output spaces of the training and testing datasets are the same, i.e., $\mathcal{X}^{\text{train}}=\mathcal{X}^{\text{test}}$ and $\mathcal{Y}^{\text{train}}=\mathcal{Y}^{\text{test}}$, but their distributions differ, i.e., $\mathbb{P}^{\text{train}}(X,Y)\neq\mathbb{P}^{\text{test}}(X,Y)$. The objective of OOD generalization is to train a predictor $f_{\theta}(X)$ based on $\mathcal{D}^{\text{train}}$ such that the error on $\mathbb{P}^{\text{test}}(X,Y)$ is minimized. The general mathematical formulation of OOD generalization aims to minimize the expected test loss over the unseen distribution, given that only training samples are accessible during model training. Given that $(X,Y)\sim\mathbb{P}_{\text{train}}(X,Y)\text{ during training}$,
$\min_{\theta}\mathbb{E}_{(X,Y)\sim\mathbb{P}_{\text{test}}}\left[\mathcal{L}(f_{\theta}(X),Y)\right].$ (2)
The OOD generalization method performs well, i.e., the predictor $f_{\theta}(X)$ learned from the training domain should perform well in all unseen domains.
II-B2 Time Series Domain
TS-OOG specifically addresses scenarios involving sequential data. At test time, the observed sequences follow a different distribution,
$\mathbb{P}_{\text{test}}(\{X_{t},Y_{t}\}_{t=T+1}^{T+\tau}),\quad\mathbb{P}_{\text{test}}\neq\mathbb{P}_{\text{train}}.$ (3)
TS-OOG aims to minimize expected loss over future observations drawn from a shifted distribution, given historical data for prediction. Given $\{X_{t},Y_{t}\}\sim\mathbb{P}_{\text{train}}$ during training,
$\min_{\theta}\mathbb{E}_{\{X_{t},Y_{t}\}\sim\mathbb{P}_{\text{test}}}\left[\sum_{t=T+1}^{T+\tau}\mathcal{L}(f_{\theta}(X_{t-\Delta:t-1}),Y_{t})\right],\quad$ (4)
where a forecasting model $f_{\theta}(\cdot)$ with parameters $\theta$ predicts future outcomes based on a historical data window $X_{t-\Delta:t-1}$, and $\tau$ denotes the length of the future prediction horizon.
Compared to general machine learning tasks, TS-OOG faces numerous unique challenges. Due to the sequential dependencies inherent in time series data, long-term trends and short-term fluctuations are intricately intertwined. At the same time, its non-stationary nature leads to dynamic distribution shifts over time [25, 26]. Additionally, periodic and seasonal patterns further complicate the generalization process. The instability of causal relationships can also result in shifts in lag effects as environmental conditions change, thereby increasing the difficulty of adapting to unseen environments. Moreover, data sparsity imposes significant constraints on effectively learning from rare events. Against this backdrop, the core objective of TS-OOG is to maintain stable and accurate performance despite non-stationary behaviors, distribution shifts, or structural changes. Achieving this goal necessitates incorporating dynamic adaptation mechanisms to accommodate unique data distributions. Additionally, robust representation learning strategies should be developed.
II-C Out-of-Distribution Generalization Bound
The generalization bound measures the predictive performance of a learning algorithm and prevents overfitting, ensuring that the model’s performance on unseen data is close to its performance on the training data [27, 28]. It provides an upper bound for the true risk with high probability, depending on the sample size and the complexity of the hypothesis space. It achieves a balance between empirical error and true error [29, 30].
Let $A$ be a learning algorithm, $\mathbb{P}$ be an unknown distribution on $\mathcal{X}\times\mathcal{Y}$, and $\delta>0$. Then, with at least a probability of $1-\delta$, for a sample $z=(z_{1},\cdots,z_{n})$ drawn from $\mathbb{P}^{n}$, the risk $L_{\mathbb{P}}(h)$ of the hypothesis $h=A(z)$ will not exceed a certain value $\hat{L}_{z}(h)+\epsilon$. That is, the general form of the generalization bound is:
$P_{n}\left(L_{\mathbb{P}}(h)\leq\hat{L}_{z}(h)+\epsilon\right)\geq 1-\delta,$ (5)
where $P_{n}$ represents the probability computed based on a sample of size $n$, $L_{\mathbb{P}}(h):=\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell(h(x),y)]$ denotes the loss of the model $h$ under the target distribution, $\hat{L}_{z}(h):=\frac{1}{n}\sum_{i=1}^{n}\ell(h(x_{i}),y_{i})$ is the empirical risk of the model $h$ over the training set $z$, and $\delta$ represents the confidence parameter.
(OOD Error Bounds in Time Series) Data samples are typically non-IID in time series modeling due to inherent temporal dependencies [31, 32]. Consequently, the classical IID assumption no longer strictly holds, necessitating the introduction of mixing process assumptions (such as $\beta$-mixing or $\phi$-mixing) to characterize sequence dependencies and extend generalization error bounds. A mixing process roughly requires that as the time gap increases, the dependence between samples gradually diminishes, approaching independence [33]. Under this premise, learning theory results from IID settings can be generalized to time series data. For instance, Kuznetsov and Mohri [34] established the first generalization bounds for time series forecasting under non-stationary mixing processes. They demonstrated that for $\beta$-mixing and $\phi$-mixing time series, generalization bounds can be derived based on Rademacher complexity [35, 36], where the upper bound incorporates the mixing coefficients and a measure of distributional discrepancy between the training and target distributions. When the statistical properties of a time series evolve (i.e., it is non-stationary), the training distribution may differ from the future test distribution, leading to an OOD scenario. Their findings indicate that even under non-IID data conditions, as long as certain mixing conditions hold, it is still possible to derive non-asymptotic generalization error bounds. Specifically, they established the following upper bounds:
For average-path generalization [33, 37, 38]
The expected generalization error over a time series of length $T$ satisfies the following upper bound.
$\begin{split}L_{\mathbb{P}}(h)\leq\hat{L}_{z}(h)+\Psi(&\beta\text{-mixing coefficient},\\ &\text{distributional discrepancy},m),\end{split}$ (6)
where $\Psi(\cdot)$ is a function dependent on several factors, including the $\beta$-mixing coefficient, which quantifies the rate at which statistical dependence between observations decreases over time, the distributional discrepancy between the training and test distributions, which accounts for potential drift, and the sample size $m$, where a larger $m$ generally results in a tighter generalization bound.
For path-dependent generalization [33, 39, 40]
Average-path generalization considers the expected generalization error over individual time steps while accounting for temporal dependencies. When considering the entire trajectory of a sequence as a single sample, they also provided complexity bounds based on $\phi$-mixing. For time series represented as a trajectory, the generalization error satisfies:
$\begin{split}L_{\mathbb{P}}(h)\leq\hat{L}_{z}(h)+C\cdot\Phi(&\phi\text{-mixingcoefficient},\\ &\text{sequence complexity},m),\end{split}$ (7)
where $C$ is a constant dependent on model complexity and hypothesis class capacity, $\Phi(\cdot)$ is a function dependent on several factors: the $\phi$-mixing coefficient, which quantifies the rate at which statistical dependencies decay across the trajectory; the sequence complexity, which reflects the inherent difficulty of modeling long-range dependencies in the time series; and the sample size $m$, where a larger $m$ generally results in a tighter bound, reducing the overall generalization error.
{forest}
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw, rounded corners,align=left, minimum width=2.5em, inner xsep=4pt, inner ysep=1pt, , where level=1text width=5.3em,fill=blue!10, where level=2text width=5em,font=,fill=pink!30, where level=3font=,yshift=0.26pt,fill=yellow!20, [Out-of-distribution
Generalization
in Time Series,fill=green!20 [Data Distribution
(Sec. III-A),text width=6.2em [Covariate Shift, text width=5em [ SBO [41] / CVAE [42] / Improved LIME [43] / SEGAL [44]/
Two-stage method [45] / TKNets [46] / CTOT [47] / RevIN [48] /
Connect Later [49] / ETSC [50] / TIMEX++ [51]. ] ] [Concept Shift, text width=5em [ DEXTER [52] / DIOSC [53]. ] ] ] [Representation
Learning (Sec. III-B),text width=7.4em [Decoupling-based,text width=6em [Multi-Structured Analysis: ITSR [54] / SCNN [55] / FSD [56] /
St-RKM [57] / $\beta$-VAE [58] / SRS [59] / MSGNet [60]. ] [Causality-Inspired: Causal-HMM [61] / CTSDG [62] /
MaSDEs [63] / LNN [64] / CCR [65] / DIOSC [53]. ] ] [Invariant-based, text width=5em [Invariant Risk Minimization: FOIL [66] / ROBUST [67] / TSModel [68] /
DoubleAdapt [69] / TSMixer [70] / Meta-Learning [71] / OOD-TV-IRM[72]. ] [Domain-Invariance: Two-stage method [45] / Connect Later [49] / TKNets [46] / Koodos [73] /
Diversify [74, 75] / OODRL [76] / In-N-Out [77] / SMORE [78] / EvoS [79]. ] ] [Ensemble-based, text width=5.2em [Homogeneous Ensembles: DEXTER [52] / Hybird-DWT [80] /
DTE [81, 82] / Deep ensembles[83]. ] [Heterogeneous Ensembles: CauSTG [84]. ] ] [Large Time-Series Models, text width=8.5em [Tuning-based: ForecastPFN [85] / TTMs [86] / ETP [87] /
JoLT [88] / TransferKF[89] / Time-LLM [90] / CALF [91] /
Moirai [92] / GTT [93] / LLM4TS [94] / ChatTime [95]. ] [Prompt-based: LLMTIME [96] / TimeGPT-1 [97] / TableTime [98] / CloudOps [99]. ] [Others: GPT4TS [100] / Lag-Llama [101] / MOMENT [102] /
Timer [103] / TimesFM [104] / Chronos [105]. ] ] ] [OOD Evaluation
(Sec. III-C),text width=6.5em [Extrinsic, text width=3.5em [Verifiable OOD Evaluation: Improved LIME [43] / SEGAL [44] /
TIMEX++ [51] / CauSTG [84] / SCNN [55]/ MSGNet [60]. ] [Speculative OOD Evaluation: SBO [41] / In-N-Out [77] / SMORE [78] /
PE-Att [106] / Two-stage method [45] / ACGAN [107] / LNN [64]. ] ] [Speculative, text width=3.5em [ OODRL [76] / DIVERSIFY [74, 75] / CCR [65] / Koodos [73] / FAD [108] / FOIL [66] /
CTSDG [62] / St-RKM [57] / Connect Later [49] / EvoS [79] / DIOSC [53]. ] ] ] ] ]
Figure 4: A taxonomy of methods for TS-OOG, categorized into data distribution, representation learning, and OOD evaluation, with various approaches such as covariate shift adaptation, invariant learning, ensemble methods, and LLM-based techniques.
In essence, in TS-OOG settings, if the sequence exhibits stationarity or asymptotic stationarity (i.e., satisfies mixing conditions) [109, 110], the training set can be regarded as originating from a closely related distribution, thereby theoretically ensuring bounded generalization error for future sequences subject to mild distribution shifts [111]. However, suppose a time series undergoes abrupt structural changes or distribution shifts. In that case, these shifts must be explicitly accounted for in the error bounds [112]. Similarly, McDonald et al. [113] derived non-parametric risk bounds for traditional time series models (e.g., AR [114], MA [115], and linear state-space models [116, 117, 118]), ensuring that prediction error remains controlled under weak assumptions even with finite samples. These theoretical tools provide a quantitative foundation for analyzing TS-OOG scenarios [119, 120, 121, 122].
III Taxonomy
This section will provide a detailed overview of the TS-OOG methods. As shown in Fig. 4, we discuss these methods from the following three perspectives:
  1. (1)
    Data Distribution: These methods primarily focus on addressing the differences in data distribution between the training and test datasets. Two popular approaches under this concept are: a). Covariate Shift: This assumes that the distribution of input features changes between the training and testing phases, while the conditional distribution $\mathbb{P}\left(Y\mid X\right)$ remains unchanged. b). Concept Shift: As time progresses or the environment changes, the conditional distribution $\mathbb{P}\left(Y\mid X\right)$ shifts. In other words, there is a fundamental change in the concept or environment dynamics between the training and testing stages.
  2. (2)
    Representation learning: These methods aim to reduce or eliminate the differences between data distributions by learning effective feature representations, thereby improving the model’s generalization ability. Four representative techniques are: a). Decoupling-based Methods: This approach decomposes features into task-relevant and irrelevant parts, thereby eliminating the interference of irrelevant features on the model and enhancing its robustness. b). Invariant-based Methods: This approach focuses on extracting features that remain invariant across different domains, ensuring that the model can still rely on these stable features for prediction when facing distribution shifts. c). Ensemble-based Learning: This approach combines features from multiple source domains to build a comprehensive feature representation, thereby enhancing the model’s adaptability to the target domain. d). Large Time-Series Models: Utilizing large-scale pre-trained models to capture long-term dependencies and complex patterns in time series data, thereby improving performance across different time distributions.
  3. (3)
    OOD Evaluation: These methods assess a model’s performance on OOD data and generally include two approaches: a). Extrinsic OOD Evaluation: This involves constructing or collecting OOD test datasets with real labels (or generating pseudo-labels) to directly measure the model’s prediction or classification performance across different distributions. b). Intrinsic OOD Evaluation: This involves analyzing the internal characteristics of the model (such as uncertainty quantification, consistency of feature distributions, and smoothness of responses) to indirectly infer the model’s generalization ability on OOD data.
These three methods are conceptually distinct and can be used individually or in combination to achieve better performance. We will now elaborate on each method in detail.
III-A Data Distribution
Data distribution has a decisive influence on model performance and generalization ability. However, real-world data distributions are often dynamic rather than static, frequently subject to various distribution shifts that challenge the assumptions made during training. Two common distribution shifts are covariate and concept shifts, as illustrated in Fig. 5.
III-A1 Covariate Shift
Covariate shift is a key factor affecting model generalization. To address this challenge effectively, researchers have employed advanced techniques such as generative models, data augmentation, and transfer learning, significantly enhancing robustness and interpretability. For example, a method combining soft brownian offset sampling and autoencoders can generate realistic OOD samples, reinforcing the classifier during training and improving its recognition capability on OOD data [41]. Additionally, conditional variational autoencoders have been used to generate input samples aligned with the training distribution, thereby improving the accuracy of feature importance assessments and preventing OOD issues arising from feature replacement [42]. In the domain of multivariate time series classification, an improved version of the local interpretable model-agnostic explanations approach uses generative models and adaptive weighting methods to enhance distributional consistency in the local neighborhood, thus improving the stability and credibility of explanations [43, 44, 45]. Temporal domain generalization methods, such as temporal Koopman networks and continuous-time modeling optimal transport trajectories, strengthen a model’s generalization ability in dynamic environments by aligning and modeling data distributions across temporal evolution [46, 47]. When facing different distributions, Kim et al. [48] remove non-stationary components from time series, enabling the model to maintain more stable and efficient predictive performance. Furthermore, data augmentation and transfer learning approaches, introduce targeted enhancements during fine-tuning, adapting the model to the target distribution and improving cross-domain generalization [49]. Meng et al. [50] employed meaningful perturbations and optimization techniques to interpret time series classifiers, reducing the influence of OOD data and ensuring the reliability of feature importance assessments. Meanwhile, Liu et al. [51] proposed the TimeX++ framework, which combines the information bottleneck principle and continuous-time modeling to effectively address trivial solutions and distribution bias in time series data, thereby significantly improving explanation quality.
Figure 5: Each circle represents a data instance, with different colors indicating the class to which the instance belongs. (a). Original data: The distribution of instances and the classification boundary (dashed line) remain stable. (b). Covariate shift: The distribution of instances changes (the feature distribution $\mathbb{P}(X)$ changes), but the class decision boundary $\mathbb{P}(Y|X)$ remains unchanged. (c). Concept shift: The class decision boundary $\mathbb{P}(Y|X)$ changes, but the distribution of instances (the feature distribution $\mathbb{P}(X)$) remains unchanged.
III-A2 Concept Shift
Concept shift involves a shift in the relationship between the output labels and the input features [123]. For example, in reinforcement learning, environmental dynamics may change, causing previously learned policies to perform poorly in new environments [52]. Similarly, in time series analysis, changes in inter-attribute correlations can affect a model’s generalization performance [53]. However, research specifically targeting concept shifts in time series remains limited. Concept shift places higher demands on a model’s generalization ability because the model must adapt to and recognize these semantic-level changes to maintain robustness and accuracy on OOD data. Addressing concept shift typically requires more sophisticated detection and adaptation mechanisms, such as contrastive learning frameworks and ensemble learning techniques, to enhance model robustness and adaptability in dynamically changing environments.
III-B Representation Learning
In time series data, distribution shifts due to seasonality, trends, or unexpected events often cause mismatches between training and test data. Representation learning is crucial for TS-OOG as it extracts robust features to enhance generalization. Tables II and III summarize the key characteristics of different approaches.
III-B1 Decoupling-based Methods
Decoupling-based methods address distribution shifts by separating time series into relevant and irrelevant features, thereby improving model robustness and generalization across different temporal environments. This ensures performance stability. Two major strategies are introduced below: multi-structured analysis and causality-inspired approaches.
TABLE II: Comparison of decoupling-based, invariant-based, and ensemble-based classification methods.
Category Subcategory Model Data Type
Decoupling-based Multi-Structured Analysis ITSR [54] Multivariate discrete time series
Multi-Structured Analysis SCNN [55], FSD [56], St-RKM [57], SRS [59], MSGNet [60] Multivariate time series
Multi-Structured Analysis $\beta$-VAE [58] Timestamped time series
Causality-Inspired Causal-HMM [61],CTSDG [62],LNN [64] Multivariate time series
Causality-Inspired MaSDEs [63] Multivariate time series with noise
Causality-Inspired CCR [65] Perturbed univariate time series
Causality-Inspired DIOSC [53] Cross-domain time series
Invariant-based Invariant Risk Minimization FOIL [66],TSMixer [70],Meta-Learning [71] Multivariate time series
Invariant Risk Minimization ROBUST [67],DoubleAdapt [69],OOD-TV-IRM [72] Multivariate discrete time series
Invariant Risk Minimization TSModel [68] Cross-domain time series
Domain-Invariance Two-stage [45],TKNets [46],Diversify [74],In-N-Out [77] Multivariate discrete time series
Domain-Invariance Connect Later [49] Multivariate time series
Domain-Invariance Koodos [73], OODRL [76] Continuous time series
Domain-Invariance SMORE [78], EvoS [79] Discrete time series
Ensemble-based Homogeneous Ensembles DEXTER [52], Hybird-DWT [80], DTE [81],Deep ensembles [83] Multivariate discrete time series
Heterogeneous Ensembles CauSTG [84] Multivariate discrete time series
Multi-Structured Analysis
Multi-structured analysis methods based on decoupling decompose time series data (e.g., through label decomposition or component decomposition) into multiple independent and interpretable parts such as short-term, long-term, and seasonal components. For example, ITSR [54] applies a learnable orthogonal decomposition to partition time series into invariant and correlated features, ensuring consistency of core patterns across different domains while mitigating domain shifts. SCNN [55] decomposes multivariate time series into structured and heterogeneous components, with dedicated sub-networks modeling each component to predict complex spatiotemporal patterns efficiently. In addition, FSD [56] leverages conditional distribution tests to locate features where shifts occur, thus enhancing robustness in high-dimensional and noisy environments. St-RKM [57] introduces an energy-based approach for unsupervised OOD detection, further improving generalization in multivariate data. $\beta$-VAE [58] uses latent-space modeling for efficient OOD detection, sensitively capturing changes in image features through partial decomposition of latent variables, thereby identifying OOD samples. SRS [59] detects anomalous samples by decomposing input data and using deep generative models to estimate conditional probability distributions, computing a “seasonality ratio score” particularly suitable for OOD detection in time series. Finally, MSGNet [60] employs frequency-domain analysis to decompose time series into multi-scale components and combines adaptive graph convolutions to capture cross-series correlations at different temporal scales, thus boosting both the accuracy of multivariate time series forecasting and generalization to OOD samples.
Causality-Inspired Models
The decoupling-based causal approaches mitigate spurious correlations in time series by isolating and extracting causal features, thereby significantly enhancing the model’s generalization performance on unseen distributions. Progressively moving from raw causal structures to extracting causal features subject to less local distribution shift, then to extracting causal features under more substantial global distribution shifts, these methods gradually enhance adaptability to complex time series. For instance, Causal-HMM [61] divides hidden variables into disease-related and spurious components, establishing a foundation for robust prediction by isolating causal features. Building on this, CTSDG [62] introduces structural causal models combined with cyclic latent variable models, further strengthening the modeling of causal relationships in complex time series. To handle uncertainty and noise in time series, MaSDEs [63] adopt a neural stochastic differential game framework, employing multi-agent collaboration to achieve more robust causal feature extraction. LNN [64] takes advantage of liquid neural networks’ dynamic causal modeling capabilities, improving adaptability and robustness in dynamic environments. As methods evolve, CCR [65] focuses on extracting causal features related to external disturbances through causal representation learning, enhancing predictive power under exogenous conditions. Finally, DIOSC [53] incorporates contrastive learning to optimize the causal representation of time series under independence constraints, ensuring strong model generalization across diverse scenarios.
III-B2 Invariant-based Methods
Invariant-based methods aim to extract feature representations that remain stable across different environments in dynamic time-series data, thereby addressing the challenges posed by distribution shifts. These approaches typically follow two complementary paradigms: one constrains the model to rely on invariant features via an optimized objective function (e.g., Invariant Risk Minimization), while the other eliminates domain discrepancies in the feature space (e.g., domain-invariance learning).
Invariant Risk Minimization
The core idea of Invariant Risk Minimization (IRM) is to discover feature representations that remain invariant across all training environments, thereby reducing the model’s reliance on spurious correlations specific to particular environments. For instance, FOIL [66] proposes a model-agnostic solution for invariant learning in time-series forecasting through a surrogate loss function and an environment inference module. A game-theoretic robust forecasting method [67] applies the IRM concept by simulating adversarial perturbations to historical data. Moreover, Zhang et al. [68] combine a specialized model architecture with IRM to capture cross-domain invariant features, further boosting generalization performance in time series. Wang et al. [72] leverage the total variation model and the primal-dual optimization method to enhance the model’s generalization ability in OOD scenarios. By learning consistent features across diverse environments, IRM helps models maintain robust performance when facing unseen periods or sudden events. In time-series applications, combining IRM methods with temporal structure modeling, dynamic invariant feature learning, meta-learning, adversarial training, and generative models have become a powerful toolkit for improving forecasting accuracy and generalization [69, 70, 71].
Domain-Invariance Models
Domain-invariance methods seek to construct universal feature representations that enable models to maintain stable performance across different domains or environments, thereby mitigating performance degradation caused by distribution shifts. First, they optimize the feature representations to maintain consistency across domains, effectively tackling distribution shifts [45, 49]. For example, conditional generative adversarial networks can discretize the output space and learn conditionally domain-invariant representations, improving adaptation to different interactive environments [45]. Furthermore, techniques such as the Koopman operator can map complex nonlinear temporal dynamics to linear spaces, bolstering model adaptability and stability in changing environments [46]. For example, Koodos [73] incorporates the Koopman operator to linearize time evolution, combined with a multi-scale attention mechanism to capture feature distribution evolution. This approach enables models to remain stable and adaptable over continuous temporal changes, effectively handling distribution drifts in dynamic environments. Additionally, some methods maximize distributional diversity to preserve heterogeneity while minimizing differences to reinforce domain invariance [74]. Lu et al. [75] propose an iterative framework that identifies “worst-case” potential distribution scenarios and reduces discrepancies. Furthermore, targeted enhancement strategies and adversarial training introduced during the fine-tuning phase improve the robustness of the model and the performance of the cross-domain, ensuring efficiency on unseen distributions [49, 76, 77]. Applying hyperdimensional computing and multi-scale attention mechanisms enhances generalization when dealing with multi-sensor data and continuous distribution drifts [78, 79].
III-B3 Ensemble-based Methods
We categorize existing ensemble-learning literature into two broad classes based on different learning principles: homogeneous and heterogeneous ensemble learning, as illustrated in Fig. 6.
Figure 6: Taxonomy of ensemble methods for time series data. (a). Homogeneous Ensembles: Comprising multiple models with identical structures that process input data. These models are integrated through a single aggregator to produce a unified output. (b). Heterogeneous Ensembles: Consisting of multiple models with different structures or varying hyperparameters. Each model generates its output independently, combining these outputs to form the final result.
Homogeneous Ensembles
In inhomogeneous ensemble learning, multiple weak learners with the same structure and learning principles but differing in training data or parameter initialization are combined to effectively reduce bias and variance. First, to address the common concept shift problem in reinforcement learning, integrating feature extraction and detection algorithms can improve anomaly detection accuracy. For instance, DEXTER [52] represents time series data as high-dimensional features and employs the isolation forest algorithm for ensembling, achieving remarkable anomaly detection results. Second, a key factor in enhancing the generalization performance of time-series data is strengthening the model’s ability to capture complex temporal patterns. Discrete wavelet transform can decompose raw time-series signals into multi-level features, which are then trained with deep sequence models. This approach significantly narrows the performance gap between ID and OOD samples [80]. In addition, multi-scale modeling through ensemble strategies substantially improves the model’s sensitivity to distribution shifts. For example, deep temporal ensemble frameworks extract features over varying time windows and integrate multiple models, effectively capturing the multi-scale characteristics of time series [81, 82]. Moreover, uncertainty awareness has become an important advancement in homogeneous ensembles to enhance model trustworthiness and safety in high-risk scenarios. For instance, uncertainty-aware deep ensembles [83] can quantify predictive uncertainty, issuing alerts for potential OOD samples. This capability significantly improves system safety and reliability in applications such as machinery fault diagnostics.
Heterogeneous Ensembles
Unlike homogeneous ensembles, heterogeneous ensemble learning integrates weak learners of various types or based on different learning principles, leveraging their complementary strengths in feature extraction and generalization. In dynamic environments, causal modeling provides stronger generalization capabilities for heterogeneous ensembles to handle continuously evolving complex distributions. For instance, CauSTG [84] partitions time environments and introduces a hierarchical invariance extraction module, significantly improving OOD samples’ detection and adaptation performance in dynamic environments—a valuable contribution to handling OOD issues in complex dynamic data. As a result, heterogeneous ensembles can more flexibly utilize the advantages of different models when facing uncertainty and complex distributions, exhibiting significant potential for broader applications in reinforcement learning, time-series analysis, and beyond.
III-B4 Large Time-Series Models
Time series modeling centered on LLMs leverages their extensive knowledge base and robust generalization capabilities [124]. In the following discussion, we focus on three directions, as illustrated in Fig. 7.
TABLE III: Comparison between three classification methods for large-scale modeling of time series. Missing Support refers to the ability to handle missing data by imputing or preprocessing it before feeding it into a model. Flexible Distribution is a qualitative classification rather than a quantitative measure, focusing on probabilistic predictions rather than just point estimates. “-” indicates not known.
Categories Model Public Open-source Transformer Model OOD/Zero-Shot Missing Support Flexible Distribution Trainable Token Trainable Parameter
Tuning-based ForecastPFN [85] NeurIPS ’23 Encoder-only - -
Jolt [125] NeurIPS ’23 Encoder-only - -
PatchTST [126] ICLR ’23 Encoder-only 20M -
TransferKF[89] EACL ’23 Encoder-only - -
ETP [87] ICASSP ’24 Encoder-only - -
Time-LLM [90] ICLR ’24 Encoder-Decoder 7B -
Moirai [92] ICML ’24 Encoder-only 150B 300M
LLM4TS [94] NeurIPS ’24 Encoder-only - 3.4M -
TTMs [86] NeurIPS ’24 Decoder-only - 5M -
aLLM4TS [127] ICML ’24 Encoder-Decoder - -
GTT [93] CIKM ’24 Encoder-only 200M 7M
CALF [91] AAAI ’25 Encoder-Decoder - -
ChatTime [95] AAAI ’25 Encoder-only 1B 350M
Prompt-based LLMTIME [96] NeurIPS ’23 Encoder-only - -
CloudOps [99] ArXiv ’23 Encoder-Decoder - 85.1M
TimeGPT-1 [97] ArXiv ’23 Encoder-Decoder 100B -
TableTime [98] ArXiv ’24 Encoder-only - -
Full-training-based (training from scratch) GPT4TS [100] NeurIPS ’23 Encoder-only - 84M
Lag-Llama [101] NeurIPS ’23 Decoder-only - 3M
MOMENT [102] ICML ’24 Encoder-only 100B 300M
Timer [103] ICML ’24 Decoder-only 50B 50M
TimesFM [104] ICML ’24 Decoder-only 3T 200M
Chronos [105] TMLR ’24 Encoder-Decoder 25B 700M
Tuning-based Methods
By leveraging tuning-based methods, the model retains its general representation capabilities while being adapted to specific tasks through synthetic pre-training, cross-modal knowledge transfer, and model reprogramming. This approach ensures strong generalization performance even when encountering OOD data. For example, ForecastPFN leverages large-scale synthetic data for pre-training, facilitating zero-shot prediction on novel tasks and enhancing adaptability to diverse time series distributions [85]. Similarly, TTMs adopt a lightweight hybrid architecture in combination with large-scale pre-training to achieve computationally efficient zero-/few-shot time series forecasting while maintaining high prediction accuracy [86]. Additionally, GTT [93] is an encoder model designed for zero-shot multivariate time series forecasting, capable of achieving high-accuracy predictions on OOD data. Since LLMs are predominantly trained on textual data, effectively transferring their learned knowledge to numerical time series poses a key challenge. ETP tackles this by jointly pre-training on ECG signals and textual data, enabling the LLM to understand both text and temporal signals and thereby improving generalization to medical data [87, 88]. Moreover, TransferKF investigates using natural language models to detect cardiovascular diseases, showing that medical knowledge learned from textual data can be successfully transferred to time series signal analysis [89]. To further enhance the efficiency of LLMs for time series tasks, researchers employ model reprogramming to adjust input formats for time series forecasting without altering model parameters. For instance, Time-LLM introduces specialized input formats that convert time series data into token sequences interpretable by LLMs, thereby improving the model’s adaptability to temporal data [90]. Additionally, CALF applies cross-modal fine-tuning to connect numeric and textual data, improving LLMs’ generalization capabilities for time series [91]. Unified modeling approaches for time series tasks have been proposed in pursuit of further performance improvements. Moirai employs a unified training framework, enabling LLMs to share knowledge across different types of time series data and thus enhance generalization [92]. Meanwhile, LLM4TS explores aligning data distributions from multiple domains to reinforce the robustness of LLMs in time series forecasting [94]. Building upon these advancements, ChatTime proposes a unified multi-modal time series foundation model, allowing LLMs to handle a wider range of forecasting tasks [95].
Figure 7: Different approaches to handling time series data using LLMs. (a). Tuning-based methods: Time series are processed through patches or tokenizers, producing outputs in temporal or textual formats. (b). Prompt-based methods: Custom tokenizers or prompt managers handle time series directly without requiring model tuning. (c). Full-training-based (training from scratch): Integrated with foundational models, LLMs are applied across diverse fields such as energy, finance, healthcare, weather, and transportation.
Prompt-based Methods
Prompt-based methods efficiently adapt to new distributions while keeping the overall structure of the model unchanged [128]. The central idea is to leverage the strong transfer capability that LLMs develop through natural language tasks by reformulating time-series forecasting, classification, and other problems into formats that LLMs can directly process. This approach thereby enhances OOD generalization. For instance, LLMTIME [96] can perform time-series forecasting solely through carefully crafted textual prompts, requiring no specialized training on time-series data, and it demonstrates competitive zero-shot generalization performance. Likewise, TimeGPT-1 [97] trains on time-series data from multiple domains, conferring robust transfer capabilities across various tasks and data distributions. Task reformulation is a key strategy for improving LLMs’ capacity to handle time-series data. TableTime [98] reframes time-series classification as a table-understanding task, allowing it to exploit existing table-processing techniques. This approach avoids the extensive labeled data typically required by conventional time-series classification methods and significantly boosts generalization to diverse data distributions. Meanwhile, CloudOps [99] employs large-scale pre-training on time series data related to cloud computing operations, enabling robust predictions under OOD circumstances such as sudden anomalies or system upgrades without additional fine-tuning.
Full-training-based (training from scratch)
As LLMs continue to expand their applications in time series analysis [129], researchers have proposed foundation models for time series that enhance the OOD adaptability of LLMs through large-scale pre-training, task-agnostic modeling, and cross-domain generalization. Its powerful generalization capability is precisely attributed to the implicit data augmentation effect brought by large-scale pre-training. GPT4TS [100] investigates how to leverage pre-trained language models to build a universal framework for time series analysis, allowing different types of time series tasks to be addressed without domain-specific fine-tuning and demonstrating strong cross-task generalization capabilities. Lag-Llama [101] extends time series modeling by introducing an LLM variant that integrates autoregressive and sequence-modeling abilities, thereby handling long-term dependencies in forecasting tasks and improving generalization to unseen data distributions. Building on this foundation, MOMENT [102] pre-trains on multi-domain data, enabling the model to adapt to various tasks such as prediction, classification, and anomaly detection. Meanwhile, Timer [103] explores the use of GPT-style transformers in processing time series data, highlighting the potential of these models for zero-shot forecasting. TimesFM [104] proposes a decoder-only architecture optimized for computational efficiency in time series forecasting while enhancing adaptability to non-stationary series. Finally, Chronos [105] advances the concept of a ’time series language’, where universal temporal patterns are learned so that LLMs can interpret and generalize across diverse domains.
III-C OOD Evaluation
OOD evaluation aims to quantify a model’s generalization ability under unknown data distributions. Based on the completeness of the evaluation data, existing approaches can be divided into two categories: extrinsic evaluation and intrinsic evaluation. We summarize their differences from OOD assessment in Table IV. These two complementary approaches form a comprehensive evaluation framework spanning data to model.
TABLE IV: Comparison of Extrinsic and Intrinsic OOD Evaluation.
Dimension Extrinsic OOD Evaluation Intrinsic OOD Evaluation
Verifiable Speculative
OOD test data
Label availability -
Evaluation basis True labels Uncertainty Internal features
TABLE V: Extrinsic OOD Evaluation Type.
Types Methods Datasets Types Methods Datasets
Verifiable SLIME-MTS [43] UEA Speculative SBO [41] Generating unlabeled data
Verifiable SEGAL [44] UEA Speculative In-N-Out [77] Generating pseudo-labels
Verifiable Causal-HMM [61] In-house data on PPA Speculative SMORE [78] Unlabeled USC-HAD
Verifiable TimeX++ [51] UCR Speculative PE-Att [106] Generating simulated data
Verifiable CauSTG [82] SIP; METR-LA; KnowAir; Electricity Speculative LNN [64] Generating data
Verifiable SCNN [55] Traffic; Solar-energy; Electricity; Exchange-rate Speculative VRStrokeOOD [45] Generating unlabeled data
Verifiable MSGNet [60] Flight; Weather; ETT; Exchange-Rate; Electricity Speculative ACGAN [105] Unknown attack data
III-C1 Extrinsic OOD Evaluation
Extrinsic evaluation constructs or gathers explicit OOD test data to assess the model’s generalization performance directly. According to the source and labeling status of the test data, extrinsic evaluation can be further subdivided into verifiable and speculative assessments. Table V provides a detailed breakdown of the categories of Extrinsic OOD Evaluation and their corresponding datasets.
Verifiable OOD Evaluation
Verifiable OOD evaluation measures a model’s OOD generalization performance on labeled test datasets across different time points. Although assessing models on OOD datasets seems intuitive, ensuring reliable and reproducible evaluation outcomes requires the design of rigorous protocols. These protocols should specify how distribution shifts are induced, how data environments are partitioned, and how fairness is maintained under different testing conditions. For instance, Meng et al. [43, 44] employ stable explanation methods to validate model performance in various data environments and provide quantitative metrics and visualizations, enhancing transparency and verifiability in the evaluation process. In addition, some researchers [61, 51] introduce causal modeling and information bottleneck strategies, which help reveal how latent factors affect a model’s OOD performance, thereby adding interpretability to the evaluation. Furthermore, drawing on research on capturing spatiotemporal invariance and multi-scale correlations [84, 55, 60] offers deeper insights into how models behave under real-world distribution shifts.
Speculative OOD Evaluation
In real-world scenarios, obtaining labeled OOD test data is often highly challenging. Consequently, researchers typically rely on unlabeled test data to assess model applicability or identify the optimal model. In this context, speculative OOD evaluation emerges as an effective strategy. Various methods have been explored and adopted to mitigate the lack of labels for OOD test data, including synthetic data generation [41], pseudo-label inference [77], and dynamic test customization at the inference stage [78]. Synthetic data is a common OOD evaluation strategy to simulate model behavior under complex conditions. For instance, Narwariya et al. [106] leverage the PowerTAC simulator to generate electricity consumption data under different pricing strategies, thus examining model robustness within dynamic market environments. Similarly, Yuan et al. [45] employ domain-mapping techniques to convert desktop-based gesture data into virtual reality gesture data, thereby alleviating the shortage of labeled VR datasets and providing a new source for OOD evaluation. When labels are unavailable, pseudo-label generation offers another effective approach for OOD evaluation. Tie et al. [77] propose the In-N-Out framework, which employs an auxiliary-information-guided self-training mechanism to produce pseudo-labels and optimize model adaptability. In a related vein, Zhao et al. [107] utilize auxiliary classifier generative adversarial networks to generate latent feature representations for unknown network attack types, thereby enhancing the model’s capability to identify classes it has not previously encountered. Dynamic inference strategies furnish real-time adaptability solutions for OOD evaluation. For example, Wang et al. [78] develop the SMORE framework, which leverages hyperdimensional domain adaptation at inference time to adjust a multi-sensor time series classification model dynamically. The model can quickly adapt to sensor distribution shifts during testing by exploiting a similarity-based feature reorganization mechanism. Furthermore, liquid neural networks [64], which can modify their parameters in real-time, complement generative approaches to form a complete pipeline.
III-C2 Intrinsic OOD Evaluation
Intrinsic OOD evaluation refers to inferring a model’s potential generalization capability by analyzing its intrinsic properties when test datasets are unavailable. Current research has extensively investigated critical characteristics, including distributional robustness [74, 76, 65], smoothness [73, 108], and invariance [66, 62], which fundamentally govern model generalization and enhance performance on unseen data. Distributional robustness ensures stable and reliable predictions under distribution shifts. For instance, Tonin et al. [57] validated the effectiveness of distributional robustness-based methods through similarity metrics and energy function perspectives, respectively. Meanwhile, Lu et al. [75] established a unified framework to directly quantify model responses to OOD data, while [49] demonstrated enhanced adaptability to distribution shifts via targeted data augmentation. For example, [73] proposed continuous temporal domain generalization by capturing the smooth temporal evolution of model outputs to predict their performance on unseen data, whereas [79] further explored how normalization strategies achieve sustained smooth adaptation against temporal drift. Invariance-oriented approaches focus on extracting core features insensitive to external disturbances, thereby maintaining high performance under OOD scenarios. Specifically, [62] and [65] employed causal inference and causal representation learning to disentangle confounding factors, while [53] utilized contrastive learning for feature decoupling to reinforce invariant internal representations. Notably, although more challenging in the absence of test data, this intrinsic evaluation paradigm provides deeper insights into the essence of OOD generalization. It compels models to concentrate on relevant features while maintaining stable and accurate predictions across diverse OOD scenarios.
IV APPLICATIONS
This section explores applications of TS-OOG across multiple domains, highlighting the urgent need for high generalization capacity in each, as shown in Fig. 8. As data collection environments become increasingly complex, data from various domains often exhibit distribution shifts due to external factors. Consequently, developing robust and adaptable models across domains has become paramount.
{forest}
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, rectangle, draw, rounded corners,align=left, minimum width=2.5em, inner xsep=4pt, inner ysep=1pt, , where level=1text width=5em,fill=blue!10, where level=2text width=5em,font=,fill=pink!30, where level=3font=,yshift=0.26pt,fill=yellow!20, [TS-OOG
Applications,fill=green!20 [Transportation,text width=5.5em [Autonomous driving,text width=6.5em [e.g. LNNs [64]/ [130] / [131]. ] ] [Smart city,text width=4em [e.g. TSModel [68] / [132] / [133]. ] ] ] [Environment,text width=5em [Air quality,text width=4em [e.g. DAQFF [134]. ] ] [Remote sensing,text width=6em [e.g. AQNet [135]. ] ] [Energy consumption,text width=7em [e.g. PE-Att [106] / [136]. ] ] ] [Public health,text width=5em [Drug discovery,text width=5em [e.g. EvoS [79]. ] ] [Mental health,text width=4.5em [e.g. EEG-ECG [137]. ] ] ] [Public safety,text width=5em [Cybersecurity,text width=4.5em [e.g. ACGAN [107] / AFFAR [138]. ] ] ] [Others,text width=3em [Robust,text width=3em [e.g. FOIL [66]. ] ] [Finance,text width=3.5em [e.g. [139]. ] ] [Fault Diagnosis,text width=5em [e.g. Deep ensembles [83]. ] ] [Human activity recognition,text width=5.5em [e.g. DTE [81, 82]/ [45]. ] ] ] ]
Figure 8: Several applications of TS-OOG.
IV-1 Transportation
In the transportation, autonomous driving systems and smart city infrastructures rely on time-series data from various sensors, such as cameras and radar. Data gathered by autonomous vehicles often undergo significant distribution shifts due to weather conditions, lighting variations, and road surface changes [130, 64]. Thus, Researchers focus on real-time correction of distributional shifts to ensure correct decision-making under extreme conditions [131]. Meanwhile, the multitude of sensors used in smart cities (e.g., traffic flow monitoring, public facilities, environmental detection) generate data that, when effectively integrated via TS-OOG techniques, can substantially improve the intelligence of urban management [132, 68, 133]. In transportation, demand for data distribution–agnostic solutions extends to real-time decision-making and large-scale data integration and management.
IV-2 Environment
Air quality data frequently exhibit dynamic properties influenced by seasonal changes, pollution incidents, and meteorological fluctuations [134]. In contrast, remote sensing data suffer from inconsistencies in collection devices, angles, and geographic locations [135]. Similarly, energy consumption data are subject to multiple contributing factors—economic activity, climate variability, and policy interventions—rendering time-series patterns highly complex and volatile [106]. To capture these changing dynamics, researchers utilize TS-OOG methods to model data trends, achieving greater predictive accuracy and stability in environmental monitoring and management [136]. This underscores the importance of high generalization capacity in managing highly variable environmental data.
IV-3 Public Health
TS-OOG also proves advantageous in public health. During AI-assisted drug discovery, abundant molecular dynamics simulations, experimental data, and clinical trial results undergo distribution shifts due to varying experimental conditions and individual patient differences. TS-OOG methodologies enable models to maintain high consistency and predictive accuracy across these diverse experimental settings, expediting drug development [79]. Additionally, psychological health monitoring depends on physiological and behavioral data, which can differ considerably among different populations and contexts. Constructing diagnostic systems capable of handling distributional shifts allows earlier detection of potential mental health issues. For instance, Yang et al. [137] integrate EEG and ECG data to achieve robust generalization in epilepsy detection. In public health, addressing distributional shifts is critical for accurate diagnosis and prediction.
IV-4 Public Safety
System logs and network traffic data in cybersecurity often exhibit anomalous patterns under novel attacks. By incorporating TS-OOG techniques, models become more sensitive to these irregular signals, enabling timely identification and prevention of unknown network threats [107]. Moreover, human activity recognition systems frequently rely on temporal data collected from wearable or surveillance devices; such data are easily affected by individual variability and environmental factors, leading to distributional shifts [138]. Researchers are, therefore, actively pursuing cross-user and cross-scenario modeling methods to achieve stable recognition. These applications demonstrate how OOD generalization methods can effectively mitigate challenges posed by data changes, thereby strengthening public safety.
IV-5 Others
It is worth noting that TS-OOG techniques exhibit wide-ranging potential in other areas. In finance, for instance, stock prices, trading volumes, and macroeconomic indicators are inherently non-stationary, heavily influenced by regulatory changes, market sentiment, and external economic environments. OOD generalization can help financial models adapt to these dynamic shifts, offering more reliable data support for market prediction, risk management, and anomaly detection [139] Similarly, for industrial equipment fault diagnostics [83], OOD generalization shows promise in anomaly detection and preventive strategies. By continuously exploring and applying TS-OOG techniques across diverse domains, systems achieve higher robustness under extreme conditions and more trustworthy data-driven decision-making.
V CHALLENGES AND FUTURE DIRECTIONS
This section discusses several challenges and future directions for TS-OOG.
V-1 Dynamic Distribution Shifts and Continual Adaptation Mechanisms
In open and dynamic environments, machine learning systems should effectively handle continuous changes in data distributions, which primarily manifest as covariate shift and concept drift [140, 141]. These shifts lead to degradation in model prediction accuracy, where traditional batch training methods fail to capture gradual changes in data streams and may cause critical decision failures in high-stakes applications. Current research focuses on constructing cognitive closed-loop systems that enhance model adaptability through dynamic detection [142, 143], continual adaptation [144, 145], and trustworthiness verification [83, 146]. Future investigations should prioritize the optimization of online algorithms for efficient shift identification under constrained computational resources, establish task sequence knowledge evaluation frameworks to quantify catastrophic forgetting effects in continual learning and derive generalization error bounds in dynamic environments to provide theoretical foundations for system design. As machine learning systems progressively adapt to environmental feedback, they will achieve self-evolution and synergistic development in dynamic environments.
V-2 Constructing Robust Invariant Representations
Time series data are often influenced by seasonality, trends, and abrupt events, making it challenging to capture cross-period and cross-domain invariant information [147, 148]. Future research may focus on feature-separation strategies grounded in decoupling, causal inference, and the information bottleneck principle, thereby isolating stable patterns while adapting to distribution shifts [149, 150]. For instance, multi-scale decomposition can be used to model short-term cyclical fluctuations and long-term trends [151]. At the same time, causal graphs combined with adversarial training can suppress non-essential confounding variables; both approaches help models adapt to distribution shifts. Developing invariant representations will provide a solid theoretical and methodological foundation to enhance generalization in unfamiliar environments.
V-3 Multi-modal Unified Large-Scale Time Series Models
Existing time series models often struggle in zero-/few-shot scenarios with OOD generalization, encountering bottlenecks in cross-domain knowledge transfer, limited interpretability, and insufficient robustness [152, 153]. A promising direction is to combine large-scale multi-modal pre-training, task reformulation, and model reprogramming to construct a unified large-scale time series modeling framework [154]. Such a framework can deliver efficient generalization and in-depth analytics across finance, healthcare, and industrial domains by incorporating domain adaptation mechanisms. Developing a multi-modal unified large time series model will establish a strong technical foundation for intelligent decision-making and automated systems in various industries, thereby driving a new wave of growth in artificial intelligence applications.
V-4 Uncertainty Quantification and Benchmark OOD Evaluation for Time Series
Building a comprehensive, dynamic, and adaptive OOD evaluation system requires accurately quantifying model uncertainty while thoroughly assessing OOD performance in time series data [155, 156]. Key challenges include limited data coverage and evolving distributions. Insufficient data diversity compromises model generalization in noisy or extreme scenarios [157]. In contrast, continuous changes in data and business requirements call for real-time monitoring and ongoing optimization of evaluation procedures. Moreover, the absence of standardized, interpretable evaluation metrics hinders effective model improvement, and difficulties associated with collecting regulatory-compliant. Possible solutions include constructing layered, multi-modal datasets to enhance coverage and diversity, implementing real-time distribution monitoring and online learning for continual updates, and establishing interpretable multidimensional metrics to support diagnostic and visualization needs. Ultimately, we view the development of a unified framework that enables cross-platform and cross-model integration while combining uncertainty quantification with time series OOD evaluation as both necessary and feasible.
V-5 Explainable AI for TS-OOG
Explainability is crucial for enhancing the transparency, trustworthiness, and accountability of AI systems, particularly in high-risk areas such as healthcare [158] and autonomous driving [159]. As AI applications expand, balancing model robustness and explainability in dynamic environments has become increasingly urgent [160, 161, 162]. While deep learning models excel at capturing temporal dependencies and delivering high predictive accuracy, their decision-making processes often lack transparency in OOD generalization, undermining trust in high-stakes scenarios [163, 164, 165]. Furthermore, to ensure transparent and secure decision-making, developers and domain experts must thoroughly examine and understand such models before deployment. Therefore, future research and applications should delve into these models’ internal evolution and decision logic in dynamic time-series contexts while devising continual learning strategies that uphold performance and interpretability. We can foster the real-world implementation of explainable AI in complex and evolving environments by developing lightweight, robust time series analysis frameworks that integrate domain knowledge.
VI Conclusion
This paper provides a comprehensive review of the current state of research on TS-OOG. From data distribution and representation learning to OOD evaluation, we have summarized the fundamental ideas, technical implementations, and applications of key methodologies while examining the challenges and future directions in this area. Overall, developing robust invariant representations, unified multimodal large-scale models, and robust OOD evaluation frameworks will be instrumental in advancing this field. This survey will offer systematic theoretical guidance and practical references for researchers, facilitate the evolution of time series analysis toward more robust and interpretable next-generation systems, and ultimately enable the construction of highly generalizable time series models under complex real-world conditions.